OpenStudy (anonymous):

-5x+8y=42 Find the value for y when x=6.

OpenStudy (anonymous):

-5*6 +8y = 42 -30 +8y = 42 8y = 42+30 8y = 72 Find y

OpenStudy (anonymous):

9 ^-^

OpenStudy (anonymous):

-5x + 8y = 42 Plug in x. -5(6) + 8y = 42 -30 + 8y = 42 Now we can simply solve for y. -30 + 8y = 42 Add 30 to both sides. 8y = 72 Divide each side by 8. y = 9 <---- Answer. (: